On the slow-fast dynamical systems

Zoubir Dahmani University of Mostaganem, Algeria

In our work,we study some slow fast dynamical systems in plane, such that the reduced dynamical systems have not uniqueness of solutions in one or several points.It is to not that the usual approximation techniques can not resolve our problem that is the behavior of the solutions of the slow fast systems in the neighbourhood on the solutions of the reduced systems and we have to find others arguments to resolve the problem.